A child bounces a 50-gram superball on the sidewalk. The velocity of the superball changes from 21 m/s downward to 19 m/s upward. If the contact time with the sidewalk is 1/800 s, what is the magnitude and direction of the force exerted on the superball by the sidewalk?
Answer #1

Force = Change in momentum/Time

Change in momentum = m*(v1-v2) = 5*10^-2* (19 - (-21)) = 2

Time = 1/800

Force = 2/(1/800) = 1600N